How much pure acid should be mixed with 4 gallons of a 60% acid solution in order to get 80% acid​ solution?
Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

How much pure acid should be mixed with 4 gallons of a 60% acid solution in order to get 80% acid​ solution?
---------------
Pure is 100%
80% is the average of 60% and 100% ---> equal amounts
4 gallons

Answer by mananth(16946)


How much pure acid should be mixed with 4 gallons of a 60% acid solution in order to get 80% acid​ solution?
Pure acid 100%
Let quantity to be added be x gallons to get 80% acid
available acid 60% 4 gallons
Required 80% acid . Volume after mixing will be (4+x) gallons
100*x+60*4= (x+4) *80
100x+240 = 80x+320
20x = 80
x=4
mix 4 gallons of pure acid with the 4 gallons of the 60% acid solution to get 80% acid solution.